What is the molarity of ZnCl2 that forms when 25.0 g of zinc completely reacts with CuCl2 according to the following reaction? Assume a final volume of 265 mL . Zn(s)+CuCl2(aq)→ZnCl2(aq)+Cu(s)
Answer #1

no of moles of Zn    = W/G.A.Wt

                               = 25/65.4   = 0.382moles

volume of solution = 265ml  

                                = 0.265L

Zn(s)+CuCl2(aq)→ZnCl2(aq)+Cu(s)

1 mole of Zn react with CuCl2 to gives 1 mole of ZnCl2

0.382 moles of Zn react with CuCl2 to gives 0.382 moles of ZnCl2

molarity of ZnCl2   = no of moles/volume of solution in L

                               = 0.382/0.265   = 1.44M

molarity of ZnCl2 is 1.44M >>>>answer